OpenStudy (anonymous):

Suppose you are deteriming the growth rate of two species of plants. Species A is 12 cm tall and grows 2 cm per month. Species B is 10 cm tall and grows 3 cm per month. Which system of equations models the height of each species H(d) as a function of days d. H(d ) = 12 + 2d H(d ) = 10 + 3d H(d ) = 12 + 2d H(d ) = 3 + 10d H(d ) = 2 + 12d H(d ) = 3 + 10d H(d ) = 2 + 12d H(d ) = 10 + 3d

OpenStudy (anonymous):

Since the growth rate is constant for both species of plants, the height is a linear (not quadratic) function of the number of days since d = 0. Assuming 30 days/month: Species A: Growth rate = 2 cm/month = 2/30 = 1/15 cm/day Height on day d = 0: 12 cm Height: H(d) = 12 + d/15 Species B: Growth rate = 3 cm/month = 3/30 = 1/10 cm/day Height on day d = 0: 10 cm Height: H(d) = 10 + d/10 EDIT: None of the systems you've shown as choices match the problem since all of them have incorrect rates. The problem shows a growth rate of 2 cm per MONTH for A and 3 cm per MONTH for B. The last set of equations, H(d) = 12 + 2d and H(d) = 10+3d, would be correct if the units of "d" is months instead of days. To express d in days, you have to divide the growth rate per month by the number of days in the month.

OpenStudy (anonymous):

hmm this is the only question i dont know the answer what choice would be the best i have no idea

OpenStudy (anonymous):

to you, which if these 4 choices are most similar to my solution above?